# TabularARGN: A Flexible and Efficient Auto-Regressive Framework for Generating High-Fidelity Synthetic Data

### v1

@@ 5 Conclusion
In this paper, we introduced the Tabular Auto-Regressive Generative Network (Tabular ARGN), a novel framework for generating high-quality synthetic data for both flat and sequential tables. By leveraging a unified auto-regressive modeling approach, TabularARGN learns all conditional probabilities across features and time steps, enabling unprecedented flexibility in downstream tasks such as imputation, fairness adjustments, and group-specific conditional generation.
@@ Appendix B Heuristics and Layer Sizes
Embedding Layers $3\cdot\Big{\lceil}d_{\text{in}}^{\mathstrut 0.25}\Big{\rceil}$
Regressor Layers [16]
FlatContextCompressor [64]
HistoryCompressor [64]
Table 2: Heuristic for calculating the size of the embedding layers ($d_{\text{in}}$ is the cardinality of the input sub-column) as well as layers sizes and hidden dimensions of the components of the TabularARGN models.
@@ Appendix D Sequential Model
Both columns are treated as categorical features, ensuring compatibility with the model’s architecture. During training, these additional columns are fed to the model along with the other features, enabling the model to learn relationships between sequence positions and the observed data.
@@ Appendix F Experimental setup
Compute: We perform all experiments with flat tables on AWS p3.2xlarge instances with 8 vCPUs, 61 GiB system RAM, and one NVIDIA Tesla V100 (16 GiB of GPU memory). Experiments for sequential data are run on AWS g5.2xlarge with 8 vCPUs, 32 GiB system RAM, and one NVIDIA A10G (24 GiB of GPU memory).
Data sets: We provide all data sets used in our benchmark in a GitHub repository [^3] https://github.com/mostly-ai/paper-tabular-argn. The repository also contains the training and holdout splits used, the generated synthetic data, data pre-processing scripts necessary for the baseline methods, and scripts for running the baselines as well as TabularARGN.
Details on the flat data sets Adult (Dua & Graff, 2019) and ACS Income (Ding et al., 2021; Flood et al., 2020) sets are shown in table 3; details on the sequential data sets Baseball (Lahman, 2023) and California (Pace & Barry, 1997) in 4.
@@
Code repositories: Table 5 lists the repositories containing implementations of all methods used for the benchmarks in this paper. Throughout this benchmark, we use the default settings of TabularARGN as implemented in https://github.com/mostly-ai/mostlyai-engine. TabSyn, TabDDPM, and STaSy are all implemented in https://github.com/amazon-science/tabsyn. For implementation details of these baselines, we refer to appendix G.2 of (Zhang et al., 2024b).
@@
TabDDPM https://github.com/amazon-science/tabsyn
@@
Table 5: List of benchmark methods and their repositories.
Baselines: The baseline methods included in this study can be grouped depending on whether they process flat tables or sequential tables. Baselines in the first group include the following:
  • CT-GAN, proposed by Xu et al. (2019), is a generative model that introduces a mode-specific normalization mechanism to handle multi-modal data, and that generates numerical columns conditioned on discrete features.
  • TABDDPM (Kotelnikov et al., 2023) is a diffusion-based model that can seamlessly handle tabular data comprising both continuous and categorical features.
  • STASY (Kim et al., 2023) is a generative framework that leverages score-based generative modeling. It incorporates a self-paced learning technique and a fine-tuning strategy to stabilize denoising score matching training.
  • TABSYN, by Zhang et al. (2024b), integrates a VAE model to map mixed data types into a unified latent space, thus facilitating the application of a diffusion model to learn the distribution of the data.
The baselines for handling sequential data are the following:
  • REalTabFormer, proposed by Solatorio & Dupriez (2023), is a transformer-based framework that implements strategies for handling flat and sequential tabular data. For the former, they implement an autoregressive GPT-2 model that generates the elements of a record in a sequential fashion. For the latter, they leverage a sequence-to-sequence architecture that generates child tables conditioned on parent data.
  • SDV, introduced by Patki et al. (2016), is a framework that builds generative models of relational databases. It iterates through possible relations within a database creating a comprehensive model that enables data synthesis by sampling from any component of the database.
  • ClavaDDPM, introduced by Pang et al. (2024), employs a cluster-guided diffusion model to synthesize multi-relational databases. This method captures relationships between tables by leveraging intermediary latent representations derived from the clustering step, thus enabling the modeling of complex relational structures.
  • RC-TGAN is a generative model proposed by Gueye et al. (2023) where the generation of child table rows is conditioned on data coming from their parent rows. Moreover, this framework enables the modeling of high-order relationships in scenarios involving relational databases with multiple tables.
@@ Appendix G Further Results
This section provides a comprehensive overview of results not included in the main text. Specifically, we report generation times and split the accuracy column into univariate accuracy, bivariate accuracy, and—for sequential data—coherence. For flat table datasets, we additionally report the number of model parameters for each method (see Tab. 6).
For the Adult dataset, we include results for TabDDPM (Kotelnikov et al., 2023). While TabDDPM occasionally achieves relatively high quality, we do not report its results in the main text due to significant instability—exhibiting large variations across different runs—even for the Adult dataset. Similar to STaSy, we were unable to produce meaningful results for the ACS-Income dataset using TabDDPM.
Generation times for TabSyn are omitted, as we were unable to run inference on a GPU and had to resort to CPU execution. This makes the results incomparable to those of other methods.
Model Generation Time Univariate Acc. Bivariate Acc. Number of Parameters
Dataset: adult
TabularARGN 1 s 98.6% 97.1% 129,768
TabularARGN, DP $\epsilon=2.8$ 1 s 95.9% 91.7% 128,919
TabSyn 9 s 99.0% 97.5% 10,616,892
TabDDPM 78 s 98.0% 96.3% 11,768,942
STaSy 25 s 84.7% 76.5% 10,695,154
CTGAN 1 s 84.8% 73.2% 1,930,183
Dataset: acs-income
TabularARGN 20 s 98.9% 98.1% 465,973
TabularARGN, DP $\epsilon=2.8$ 20 s 97.6% 95.7% 465,973
TabSyn - 98.2% 96.7% 10,756,224
TabDDPM - - - -
STaSy - - - -
CTGAN - - - -
Table 6: Results for all models on the Adult and ACS-Income datasets.
For the Adult dataset, we also ran the metrics used in the TabSyn benchmark (see Tab. 7), including column-wise density and pair-wise column correlation [^4] https://docs.sdv.dev/sdmetrics, as well as the alpha precision and beta recall (Alaa et al., 2022). As the first two metrics are conceptually similar to those used in our benchmark, they yield results that align closely with the picture described in the results section 4.1.
@@
Table 7: Sample table with 5 columns.
For sequential data (see Tab. 8), we additionally report the coherence accuracy, a straightforward metric for measuring the auto-correlation within individual columns.
We also include results from SDV/PARsynthesizer (Patki et al., 2016) for the Baseball dataset. However, these results are not included in the main text, as the data quality for Baseball is very poor. Furthermore, we were unable to generate synthetic data for the California dataset due to out-of-memory errors encountered during the process. Additionally, we experimented with SDV/HMASynthesizer, but found that it does not utilize GPU acceleration and lacks parallelization. As a result, training such models is extremely time-consuming, rendering it impractical for datasets of significant size or complexity.
Model Generation Time Univariate Acc. Bivariate Acc. Coherence
Dataset: baseball
TabularARGN 39 s 95.8% 78.0% 91.5%
TabularARGN, DP $\epsilon=3.6$ 47 s 91.5% 73.0% 72.5%
SDV/PARSynthesizer 984 s 40.0% 28.0% 20.4%
REalTabFormer 1,092 s 83.6% 71.0% 79.4%
RC-TGAN 17 s 81.0% 71.0% 57.9%
CLAVA 197 s 87.9% 83.9% 66.2%
Dataset: calilfornia
TabularARGN 685 s 99.0% 98.1% 97.2%
TabularARGN, DP $\epsilon=1.2$ 682 s 96.7% 95.1% 88.4%
SDV/PARSynthesizer - - - -
REaLTabFormer 27,232 s 92.3% 89.5% 84.0%
RC-TGAN 239 s 81.0% 72.1% 66.7%
CLAVA 2,422 s 83.0% 79.0% 69.6%
Table 8: Results for all models on the Baseball and California datasets.

### v2

@@ 5 Conclusion
In this paper, we introduced the Tabular Auto-Regressive Generative Network (Tabular ARGN), a novel framework for generating high-quality synthetic data for both flat and sequential tables, and share a well-tested and actively maintained open-source implementation. Using a unified autoregressive modeling approach, TabularARGN learns all conditional probabilities across features and time steps, enabling unprecedented flexibility in downstream tasks such as imputation, fairness adjustments, and group-specific conditional generation.
@@ Appendix B Heuristics and Layer Sizes
Layer Flat Sequential Heuristic
Embeddings yes yes $3\cdot(d_{\text{in}})^{\mathstrut 0.25}$
Regressors yes yes $16\cdot\text{max}(1,\text{ln}(d_{\text{in}}))$
Context Compressor no yes $64\cdot\text{max}(1,\text{ln}(d_{\text{ctx}}))$
History Compressor no yes $32\cdot\text{max}(1,\text{ln}(d_{\text{tgt}}\cdot s_{\text{q50}}))$
Table 1: Heuristic for calculating the size of the model layers. $d_{\text{in}}$ is the cardinality of the input sub-column. $d_{\text{ctx}}$ is the full dimension of the concatenated flat context column embeddings. $d_{\text{tgt}}$ is the full dimension of the concatenated sequential target column embeddings. $s_{\text{q50}}$ is the median sequence length for the sequential table.
@@ Appendix D Sequential Model
Both columns are treated as numerical features, ensuring compatibility with the model’s architecture. During training, these additional columns are fed to the model along with the other features, enabling the model to learn relationships between sequence positions and the observed data.
@@
Max sequence window: Some datasets include sequences with very long spans, which can significantly increase training time. However, in many practical cases, the temporal correlations of interest occur over shorter scales than the full sequence length. To address this, we introduce a tunable max sequence window during the training phase of sequential models.
Training with Max Sequence Window: Instead of training on the full sequence length for each sample, a fixed window of time steps, referred to as the max sequence window, is selected. This allows the model to focus on shorter subsequences. The default max sequence window size is set to 100 time steps but can be adjusted manually to suit the characteristics of the dataset being synthesized.
For each batch, and for every sequence within that batch, a random starting point is selected for the window. This starting point is chosen uniformly, with corrections applied to ensure that time steps near the beginning and end of sequences are not underrepresented. Only the subsequence defined by this max sequence window is used for gradient calculation during training, significantly reducing computational demands for long-sequence datasets.
The generation process remains consistent with the standard approach: the sequence length is predicted in the first time step (optionally conditioned on a flat context table), and the counting index is incremented deterministically as the model generates the sequence. This ensures that the synthetic data faithfully captures the structure of the original dataset, even when the training process leverages shorter max sequence windows.
With a default value of 100, the max sequence window is only relevant for the Berka data set in this paper as the average sequence lengths of all other data sets are below 10.
@@ Appendix F Experimental setup
Compute: We perform almost all experiments on AWS g5.2xlarge instances with 8 vCPUs, 32 GiB system RAM, and one NVIDIA A10G (24 GiB of GPU memory). As a non-deep learning method, UnmaskingTrees is specifically designed to run on CPU and we opted for an AWS p3.2xlarge instance with 8 vCPUs and 61 GiB of system RAM.
Differential Privacy: Differentially private SGD is implemented using the Opacus library (Yousefpour et al., 2021). For all runs with DP-SGD, we use a noise multiplier of 1.5, a DP max. grad. norm of 1, and a $\delta=1\times 10^{-5}$.
Data sets: The Adult data set (Dua & Graff, 2019) has around 48k rows with eight categorical and six numerical features. ACS-Income (Ding et al., 2021; Flood et al., 2020) has about 1.5 million rows, containing 28 categorical and four numerical features—making it roughly 30 times larger and twice as wide as Adult, and thus a challenging real-world benchmark. Default (Yeh, 2009) contains 30,000 financial records of credit card users labeled by default status, and Shoppers (Sakar & Kastro, 2018) consists of 12,330 user sessions, capturing browsing behavior and transaction outcomes on an e-commerce platform. We initially considered an even larger data set with more columns, but all baseline methods were either too slow or required excessive GPU memory.
For sequential modeling, we present results on a two-table subset (players and fielding) of the Baseball dataset (Lahman, 2023), the California dataset (Pace & Barry, 1997) consisting of a table of households and a table of individuals, and Berka dataset (Berka, 1999), where we only consider the accounts and transactions tables. For Baseball, the players table (flat context) has about 20 thousand rows, and the sequential fielding table has 140 thousand, resulting in an average sequence length of seven. With over 600 thousand rows, the household table has about $30$ times more records than the players table, resulting in almost 1.7 million rows and an average sequence length of 2.8 in the individuals table. The accounts table has only 4,500 records, but the transactions table is notably larger, encompassing around 1 million rows. All sequence tables have a comparable number of features: 9 (Berka), 11 (Baseball) and 15 (California)
Details on the flat data sets are shown in table 2; details on the sequential data sets, in 3.
@@
Baselines: The baseline methods included in this study can be grouped depending on whether they process flat tables or sequential tables.
For flat tables, we include CT-GAN, a generative model introduced by Xu et al. (2019) that applies mode-specific normalization to handle multi-modal data while generating numerical features conditioned on categorical ones. STaSy (Kim et al., 2023) builds on score-based generative modeling, incorporating self-paced learning and fine-tuning strategies to improve the stability of denoising score matching. TabSyn (Zhang et al., 2024b) adopts a VAE to project mixed data types into a common latent space, allowing a diffusion model to better capture the underlying data distribution. Unmasking Trees, proposed by McCarter (2024), is an alternative route to deep learning models. This framework leverages gradient-boosted decision trees to incrementally unmask individual features.
For sequential tables, we consider REalTabFormer (Solatorio & Dupriez, 2023), a transformer-based framework that generates flat records autoregressively using a GPT-2 model while employing a sequence-to-sequence approach to synthesize child tables conditioned on parent data. SDV (Patki et al., 2016) takes a different approach by iteratively modeling relationships across a relational database, enabling flexible data synthesis. ClavaDDPM (Pang et al., 2024) introduces a cluster-guided diffusion mechanism that captures inter-table dependencies via latent representations, improving the modeling of complex multi-relational structures. Finally, RC-TGAN (Gueye et al., 2023) conditions the generation of child table rows on parent data, extending its modeling capacity to high-order relationships across multiple interconnected tables.
Code repositories: Table 4 lists the repositories containing implementations of all methods used for the benchmarks in this paper. Throughout this benchmark, we use the default settings of TabularARGN as implemented in x. TabSyn and STaSy are implemented in https://github.com/amazon-science/tabsyn. For implementation details of these baselines, we refer to appendix G.2 of (Zhang et al., 2024b).
@@
UnmaskingTrees https://github.com/calvinmccarter/unmasking-trees
@@
Table 4: List of benchmark methods and their repositories. As of January 2025, the TabDAR repo is not publicly available. We thank (Zhang et al., 2024a) for providing their code upfront.
@@ Appendix G Further Results
This section provides a comprehensive overview of results not included in the main text. Specifically, we report results on additional data sets (see description in F) for both flat and sequential tables.
Similar to the Adult data set, TabularARGN performs on par concerning accuracy on the ACS-Income, Default, and Shoppers data set with SOTA models (see Fig. 4). For some data sets, it marginally surpasses them. Throughout all data sets, the training time of TabularARGN is the smallest, often by a large margin. The strongest competitor in the accuracy and accuracy-efficiency trade-off is TabSyn. Applying DP-SGD increases training time and reduces accuracy. The drop in accuracy is smaller for data sets with a large sample size as the signal of individual records is better protected.
The ACS-Income data set has considerably more rows (approx. 700k in the training split) than all the other ones. This not only increases training time but makes running a subset of models difficult. UnmaskingTrees caused an OOM error on a machine with 61 GiB of RAM and CTGAN did not finish after 16 hours of training. STaSy did finish but resulted in very low accuracy numbers. We refrain from publishing these numbers as they might stem from issues other than the method itself.
Figure 4: Accuracy (top row) and Training Time (bottom row) for the flat datasets Adult, ACS-Income, Default, and Shoppers and various generative models - TabularARGN, TabularARGN with DP, TabMT, TabSyn, CTGAN, STaSy, and UnmaskingTree. Reported values are averages over five full training and generation runs, error bars indicate the minimum and maximum values. As UnmaskingTrees is a non-Deep Learning approach, it is run on CPU. For comparison, Training of TabularARGN on the Adult with a CPU takes approximately 2.5 minutes.
For the Adult dataset, we also ran the metrics used in the TabSyn benchmark (see Tab. 5), including column-wise density and pair-wise column correlation [^3] https://docs.sdv.dev/sdmetrics, as well as the alpha precision and beta recall (Alaa et al., 2022). As the first two metrics are conceptually similar to those used in our benchmark, they yield results that align closely with the picture described in the results section 4.1. Alpha precision and beta recall measure the higher-order fidelity of the synthetic data.
@@
Table 5: SDV (Patki et al., 2016) and high-order metrics (Alaa et al., 2022).
For sequential data (see Fig. 5) TabularARGN achieves the highest accuracy values across all methods. In terms of runtime, it outperforms almost all of them with the only exception of ClavaDDPM for the California data set. However, ClavaDDPM lags in accuracy underpinning the strong accuracy-efficiency trade-off of TabularARGN. RC-TGAN fails on the Berka and SDV on the California data set with an OOM error.
Applying DP-SGD has the same effect on sequential as it has on flat data. The by far strongest decrease in accuracy happens for the Berka data set which has only 2250 training subjects and the longest sequences of all benchmark data sets.
Figure 5: Accuracy (top row) and Training Time (bottom row) for the sequential datasets Baseball, California, and Berka and various generative models - TabularARGN, TabularARGN with DP, ClavaDDPM, RCTGAN, RealTabFormer, and SDV/ParSynthesizer. Reported values are averages over at least three full training and generation runs, error bars indicate the minimum and maximum values. RCTGAN and SDV/ParSynthesizer fail to train on the Berka and California datasets, respectively, due to out-of-memory (OOM) issues.
We summarize the key results from Section 4 in Table 6 for flat datasets and Table 7, presenting a detailed evaluation of synthetic data quality across all relevant metrics. For TabularARGN with DP (TabularARGN-DP) we report the average privacy budgets $\epsilon$ in Table 8. The first table reports accuracy, univariate accuracy, bivariate accuracy, coherence, DCR share, and generation time. In the second table we add coherence, a metric that assesses the quality of sequential tables specifically. These metrics offer a comprehensive assessment of the models’ ability to capture statistical dependencies and structural patterns in both flat and sequential tabular data.
Metric Model Adult Default Shoppers ACS-Income
Accuracy TabMT 0.96 [0.95, 0.96] 0.93 [0.92, 0.95] 0.94 [0.91, 0.95] 0.97 [0.97, 0.97]
TabSyn 0.98 [0.96, 0.98] 0.95 [0.94, 0.96] 0.94 [0.93, 0.95] 0.95 [0.94, 0.96]
STaSy 0.87 [0.85, 0.88] 0.89 [0.85, 0.91] 0.80 [0.71, 0.86] -
CTGAN 0.77 [0.75, 0.78] 0.77 [0.76, 0.78] 0.86 [0.83, 0.88] -
Unmasking Trees 0.86 [0.83, 0.87] 0.78 [0.78, 0.78] 0.76 [0.76, 0.76] -
TabularARGN 0.98 [0.97, 0.98] 0.92 [0.91, 0.93] 0.95 [0.95, 0.95] 0.98 [0.98, 0.98]
TabularARGN-DP 0.94 [0.93, 0.95] 0.84 [0.83, 0.85] 0.85 [0.82, 0.89] 0.96 [0.96, 0.96]
Univariate Accuracy TabMT 0.97 [0.96, 0.97] 0.95 [0.94, 0.97] 0.95 [0.93, 0.97] 0.98 [0.98, 0.98]
TabSyn 0.99 [0.98, 0.99] 0.97 [0.95, 0.98] 0.96 [0.95, 0.97] 0.96 [0.95, 0.97]
STaSy 0.90 [0.88, 0.90] 0.92 [0.88, 0.94] 0.85 [0.77, 0.89] -
CTGAN 0.83 [0.82, 0.84] 0.83 [0.82, 0.84] 0.90 [0.88, 0.91] -
Unmasking Trees 0.90 [0.87, 0.91] 0.83 [0.83, 0.83] 0.82 [0.82, 0.82] -
TabularARGN 0.99 [0.98, 0.99] 0.94 [0.93, 0.95] 0.97 [0.96, 0.97] 0.99 [0.99, 0.99]
TabularARGN-DP 0.96 [0.95, 0.97] 0.91 [0.90, 0.91] 0.90 [0.87, 0.93] 0.98 [0.97, 0.98]
Bivariate Accuracy TabMT 0.94 [0.93, 0.95] 0.92 [0.90, 0.94] 0.92 [0.89, 0.94] 0.96 [0.96, 0.97]
TabSyn 0.97 [0.95, 0.98] 0.94 [0.92, 0.94] 0.93 [0.91, 0.94] 0.93 [0.92, 0.95]
STaSy 0.84 [0.82, 0.85] 0.86 [0.81, 0.89] 0.76 [0.65, 0.83] -
CTGAN 0.70 [0.68, 0.72] 0.71 [0.70, 0.72] 0.82 [0.79, 0.84] -
Unmasking Trees 0.83 [0.79, 0.84] 0.73 [0.73, 0.74] 0.71 [0.70, 0.71] -
TabularARGN 0.97 [0.96, 0.97] 0.90 [0.89, 0.91] 0.94 [0.93, 0.94] 0.98 [0.98, 0.98]
TabularARGN-DP 0.92 [0.91, 0.93] 0.77 [0.76, 0.78] 0.81 [0.77, 0.86] 0.95 [0.95, 0.96]
DCR Share TabMT 0.51 [0.51, 0.51] 0.50 [0.50, 0.50] 0.50 [0.50, 0.50] 0.50 [0.50, 0.50]
TabSyn 0.50 [0.50, 0.50] 0.50 [0.50, 0.50] 0.50 [0.50, 0.50] 0.50 [0.50, 0.50]
STaSy 0.50 [0.50, 0.50] 0.50 [0.50, 0.50] 0.50 [0.50, 0.50] -
CTGAN 0.50 [0.50, 0.50] 0.50 [0.50, 0.50] 0.50 [0.50, 0.50] -
Unmasking Trees 0.50 [0.50, 0.50] 0.50 [0.50, 0.50] 0.50 [0.50, 0.50] -
TabularARGN 0.51 [0.51, 0.51] 0.50 [0.50, 0.50] 0.50 [0.50, 0.50] 0.50 [0.50, 0.50]
TabularARGN-DP 0.50 [0.50, 0.50] 0.50 [0.50, 0.50] 0.50 [0.50, 0.50] 0.50 [0.50, 0.50]
Generation Time (s) TabMT 10.91 [10.90, 10.92] 12.41 [12.38, 12.49] 7.28 [7.26, 7.30] -
TabSyn 7.21 [7.00, 7.27] 6.33 [6.27, 6.57] 4.52 [4.52, 4.53] -
STaSy 16.83 [14.42, 21.87] 10.10 [9.86, 10.62] 8.05 [7.44, 9.74] -
CTGAN 0.15 [0.15, 0.16] 0.17 [0.16, 0.17] 0.09 [0.09, 0.10] -
Unmasking Trees 641.10 [634.63, 646.02] 663.34 [652.94, 677.04] 187.09 [181.14, 197.17] -
TabularARGN 0.36 [0.33, 0.42] 0.77 [0.76, 0.78] 0.30 [0.30, 0.31] 15.00 [15.00, 15.00]
TabularARGN-DP 0.37 [0.34, 0.43] 0.78 [0.77, 0.78] 0.31 [0.31, 0.32] 14.80 [14.67, 14.94]
Table 6: Evaluation of synthetic data quality across flat datasets, averaged over five independent runs. For each metric, we present the mean, minimum, and maximum values as Mean[Min, Max]. The highest mean accuracy for each dataset is highlighted in bold. Similarly, the most efficient generation times are highlighted. Missing generation times for the ACS-Income data set correspond to generation runs that yielded an Out-of-Memory error in GPU. We do report values for Unmasking Trees as it is designed for training and inference on CPU. Regarding DCR Share, values close to 0.5 indicate the model is not overfitted. Missing values in fields other than Generation Time correspond to experiments that yielded an Out-of-Memory error or where a single experiment ran for more than 16 hrs.
Metric Model Baseball Berka California
Accuracy ClavaDDPM 0.79 [0.79, 0.80] 0.49 [0.49, 0.49] 0.77 [0.77, 0.78]
REaLTabFormer 0.77 [0.77, 0.78] 0.74 [0.73, 0.75] 0.89 [0.88, 0.92]
RC-TGAN 0.70 [0.69, 0.71] - 0.72 [0.69, 0.74]
SDV/PARSynthesizer 0.30 [0.29, 0.30] 0.27 [0.26, 0.28] -
TabularARGN 0.88 [0.87, 0.89] 0.79 [0.77, 0.80] 0.99 [0.98, 0.99]
TabularARGN-DP 0.79 [0.78, 0.80] 0.45 [0.44, 0.47] 0.94 [0.94, 0.94]
Univariate Accuracy ClavaDDPM 0.88 [0.88, 0.88] 0.63 [0.63, 0.63] 0.83 [0.83, 0.83]
REaLTabFormer 0.83 [0.82, 0.84] 0.81 [0.80, 0.81] 0.92 [0.91, 0.93]
RC-TGAN 0.80 [0.79, 0.81] - 0.80 [0.79, 0.81]
SDV/PARSynthesizer 0.40 [0.40, 0.41] 0.34 [0.33, 0.35] -
TabularARGN 0.95 [0.94, 0.96] 0.87 [0.85, 0.88] 0.99 [0.98, 0.99]
TabularARGN-DP 0.92 [0.92, 0.93] 0.56 [0.55, 0.59] 0.97 [0.97, 0.97]
Bivariate Accuracy ClavaDDPM 0.84 [0.84, 0.85] 0.63 [0.63, 0.64] 0.80 [0.80, 0.80]
REaLTabFormer 0.70 [0.69, 0.71] 0.63 [0.62, 0.64] 0.90 [0.89, 0.91]
RC-TGAN 0.71 [0.70, 0.72] - 0.71 [0.70, 0.72]
SDV/PARSynthesizer 0.28 [0.27, 0.29] 0.23 [0.22, 0.24] -
TabularARGN 0.77 [0.76, 0.78] 0.68 [0.66, 0.69] 0.98 [0.98, 0.98]
TabularARGN-DP 0.74 [0.73, 0.74] 0.36 [0.35, 0.38] 0.95 [0.95, 0.95]
DCR Share ClavaDDPM 0.50 [0.50, 0.50] 0.50 [0.49, 0.50] 0.50 [0.50, 0.50]
REaLTabFormer 0.50 [0.50, 0.50] 0.52 [0.51, 0.53] 0.50 [0.50, 0.50]
RC-TGAN 0.49 [0.47, 0.51] - 0.50 [0.50, 0.50]
SDV/PARSynthesizer 0.50 [0.50, 0.50] 0.51 [0.51, 0.52] -
TabularARGN 0.50 [0.49, 0.50] 0.50 [0.49, 0.50] 0.50 [0.50, 0.50]
TabularARGN-DP 0.50 [0.50, 0.50] 0.50 [0.50, 0.51] 0.50 [0.50, 0.50]
Coherence ClavaDDPM 0.66 [0.66, 0.67] 0.38 [0.38, 0.39] 0.70 [0.70, 0.70]
REaLTabFormer 0.79 [0.79, 0.79] 0.76 [0.75, 0.76] 0.85 [0.84, 0.85]
RC-TGAN 0.59 [0.58, 0.61] - 0.59 [0.58, 0.61]
SDV/PARSynthesizer 0.20 [0.20, 0.20] 0.23 [0.22, 0.24] -
TabularARGN 0.91 [0.90, 0.92] 0.82 [0.81, 0.83] 0.98 [0.98, 0.98]
TabularARGN-DP 0.72 [0.72, 0.73] 0.42 [0.41, 0.43] 0.89 [0.88, 0.89]
Generation Time ClavaDDPM 196.05 [192.76, 197.87] 1299.56 [1269.92, 1319.62] 2419.92 [2408.43, 2434.85]
REaLTabFormer 1058.41 [1044.54, 1091.67] 17169.53 [15366.52, 18395.99] 23950.47 [16713.84, 27232.35]
RC-TGAN 16.69 [12.30, 21.92] - 249.29 [239.07, 259.08]
SDV/PARSynthesizer 1010.61 [984.48, 1044.59] 1162.28 [881.63, 1774.27] -
TabularARGN 48.55 [48.29, 48.99] 64.91 [64.10, 66.81] 520.21 [519.57, 521.01]
TabularARGN-DP 35.17 [35.08, 35.33] 70.30 [68.46, 71.69] 620.31 [618.81, 622.32]
Table 7: Evaluation of synthetic data quality across sequential datasets, averaged over five independent runs. For each metric, we present the mean, minimum, and maximum values as Mean[Min, Max]. The highest mean accuracy for each dataset is highlighted in bold. Similarly, the most efficient generation times are highlighted. Regarding DCR Share, values close to 0.5 indicate the model is not overfitted. Missing values correspond to experiments that yielded an Out-of-Memory error or where a single experiment ran for more than 16 hrs.
Table Type Dataset Average $\epsilon$ Min/Max $\epsilon$
Flat ACS-Income 1.670 1.44/1.78
Flat Adult 2.386 2.11/2.59
Flat Default 2.130 1.32/2.92
Flat Shoppers 2.224 1.52/2.6
Sequential Baseball 3.826 3.43/4.18
Sequential Berka 1.866 1.59/1.94
Sequential California 1.238 1.23/1.24
Table 8: Average (over five independent runs) and Min/Max $\epsilon$ values for differentially private TabularARGN-DP models across datasets. We report the $\epsilon$ values reached at model convergence.